The length of a rectangle is 2 cm longer than its width. If the perimeter of the rectangle is 68 cm, find its length and width

Answer by ikleyn(52778)

.

    w + (w+2) = 68/2


    2w        = 34 - 2 = 32

     w                 = 32/2 = 16.


ANSWER.  16 cm  (the width) and  18 cm (the length).

Solved.